### v2

# VRWKV-Editor: Reducing Quadratic Complexity in Transformer-Based Video Editing

Abstract
In light of recent progress in video editing, deep learning models focusing on both spatial and temporal dependencies have emerged as the primary method. However, these models suffer from the quadratic computational complexity of traditional attention mechanisms, making them difficult to adapt to long-duration and high-resolution videos. This limitation restricts their applicability in practical contexts such as real-time video processing. To tackle this challenge, we introduce a method to reduce both time and space complexity of these systems by proposing VRWKV-Editor, a novel video editing model that integrates a linear spatio-temporal aggregation module into video-based diffusion models. VRWKV-Editor leverages bidirectional weighted key-value recurrence mechanism of the RWKV transformer to capture global dependencies while preserving temporal coherence, achieving linear complexity without sacrificing quality. Extensive experiments demonstrate that the proposed method achieves up to 3.7× speedup and 60% lower memory usage compared to state-of-the-art diffusion-based video editing methods, while maintaining competitive performance in frame consistency and text alignment. Furthermore, a comparative analysis we conducted on videos with different sequence lengths confirms that the gap in editing speed between our approach and architectures with self-attention becomes more significant with long videos.
Figure 1: Representative results produced by VRWKV-Editor. The proposed method is capable of object replacement, background modification, and style transformation, while consistently preserving the motion dynamics and visual characteristics of the original video. Additional qualitative examples are provided on the project page: VRWKV-Editor.
1 Introduction
The Transformer architecture has emerged as a dominant paradigm, revolutionizing fields from Natural Language Processing (NLP) to Computer Vision. Central to its effectiveness is the self-attention mechanism, which captures the long-range dependencies within a sequence. However, a significant limitation of traditional attention is its computational and memory complexity, which scales quadratically with the input sequence length. This presents a critical bottleneck for applications with limited computational resources involving high-resolution data or real-time processing constraints.
In visual tasks, Vision Transformers (ViTs) [1] and their extensions [2, 3], while demonstrating remarkable performance in image understanding, also inherit this problem and suffer from quadratic computational complexity. Videos are characterized by strong spatiotemporal dependencies, which distinguishes them from collections of independent images. Consequently, models must ensure consistency within and across frames. This dual requirement leads to an additional increase in model’s complexity.
Within the framework of text-driven video editing, optimized architectures such as Swin-transformer [2] also presents the quadratic computational bottleneck that limits scalability despite its effectiveness in handling local spatial relationships. Reducing this complexity is thus critical for both research and industry, as efficient models lower costs and enable real-time deployment in consumer applications such as mobile video editors, while also allowing researchers to experiment and iterate without relying on large-scale computational resources.
Concurrently, modern RNN-like architectures in NLP have emerged as efficient solutions for processing lengthy sequences linearly while preserving the parallelization advantages of transformers. Illustratively, RWKV, Mamba, xLSTM, and Fast Transformer [4, 5, 6, 7] each employ a distinct mechanism to linearize the classical Transformer. However, these methods are fundamentally based on two main ideas: either using an aggregation mechanism with a linear kernel in place of the softmax function, or replacing the dot-product with element-wise multiplication as a means of linearization.
In accordance with this progress, VRWKV [8] extends RWKV two-dimensional vision tasks such as image classification, detection and segmentation, demonstrating robust global processing capabilities while efficiently handling sparse inputs and maintaining linear computational complexity. Notably, VRWKV provides reduced spatial aggregation complexity and enables seamless processing of high-resolution images without requiring window operations.
Expanding on these insights, we present a new video editing framework, termed VRWKV-Editor, which leverages the linear complexity and global attention capabilities of this mechanism. Furthermore, rather than employing straightforward concatenation in the skip connections of U-Net architecture, we utilize bidirectional processing inspired by VRWKV’s Bi-WKV mechanism to merge feature maps from corresponding encoding paths with those from preceding decoding up-convolutional layers. This design enables the model to learn continuous motion patterns in input videos and transfer them to edited outputs according to user prompts, while significantly reducing computational complexity. Remarkably, our experimental results demonstrate that this system achieves computational efficiency superior to that of state-of-the-art methods, all while retaining comparable performance. In general, this work presents the following contributions:
    We propose VRWKV-Editor, a novel video editing framework that offers a cost-effective and efficient solution with reduced computational requirements. It preserves the ability to capture long-range dependencies and flexibly handle sparse inputs, while reducing complexity to a linear scale. This reduction removes the need for window operations when processing long video sequences, making VRWKV-Editor a more scalable and practical approach for video editing tasks.
    We demonstrate significant improvements in computational efficiency through the integration of VRWKV’s bidirectional attention mechanism, reducing model complexity without performance degradation.
    We provide comprehensive experimental validation showing that linear attention mechanisms can effectively replace traditional quadratic attention in video understanding tasks while offering substantial computational advantages.
The remainder of this paper is organized as follows: Section 2 reviews related work in video editing and linear attention mechanisms. Sections 3 and 4 detail background concepts relevant to our study and the proposed VRWKV-Editor architecture with its key components. Section 5 presents a comprehensive analysis and experimental results. In Section 6, we highlight potential limitations of this work. Finally, Section 7 concludes with a discussion of the study’s implications and potential directions for future research.
2 Related works
This section discusses four research areas related to our approach; video editing methods using diffusion models, which highlight the computational challenges inherent in current approaches; linear attention mechanisms, which offer efficiency gains over quadratic attention; and computational efficiency approaches in video processing, which directly motivate our integration of linear attention for scalable video editing.
2.1 Video Editing with Diffusion Models
Recent developments in diffusion model-based video editing have introduced various approaches to address the computational and quality challenges inherent in video processing [9, 10, 11, 12]. Early approaches focused on adapting image diffusion models for video tasks through temporal extensions. Tune-A-Video [11] pioneered the concept of one-shot tuning in video editing by transforming spatial self-attention layers into sparse-causal attention mechanisms. VidToMe [13] leveraged temporal redundancy through Token Merging (ToMe) to enhance consistency while reducing computational overhead. However, these methods still rely on traditional attention mechanisms with quadratic complexity, limiting their scalability to high-resolution and long-sequence video processing. Methods like Prompt-to-Prompt (P2P) [14] and Plug-and-Play (PnP) [15] have demonstrated the effectiveness of attention feature injection techniques, where cross-attention and self-attention maps from reconstruction branches guide the editing process. TokenFlow [16] and FLATTEN [17] utilize inter-frame feature similarity and optical flow-based correspondence respectively to maintain temporal consistency across video frames. While effective, these approaches face computational bottlenecks when processing high-resolution videos due to their reliance on traditional quadratic attention mechanisms. Recent works have incorporated structural guidance through depth maps, optical flow, and motion vectors to improve temporal consistency. FlowVid [18] and MoCA [19] integrate edited first frames with optical flow information to maintain motion fidelity, while approaches like ControlVideo [20] leverage multi-modal conditions for enhanced controllability.
Complementary to these editing approaches, advances in video generation architectures have explored pixel-based, latent-based, and hybrid designs [14, 21, 22]. Research has shown that pixel-based diffusion models excel in generating low-resolution videos with superior text-video alignment and motion fidelity, while latent-based models offer computational efficiency but may compromise alignment quality [22]. Hybrid approaches that combine pixel-based keyframe generation with latent-based super-resolution have demonstrated the potential to achieve both high quality and computational efficiency [23]. Effective video models incorporate temporal layers within U-Net architectures, including temporal convolution layers and temporal attention mechanisms to facilitate dynamic temporal data assimilation [24]. The integration of spatial and temporal processing remains crucial for maintaining coherent motion patterns across video sequences [25].
2.2 Linear Attention Mechanisms in Vision
The computational limitations of traditional attention mechanisms have sparked interest in linear attention alternatives that maintain performance while reducing complexity [26, 27]. Recent advances in natural language processing have introduced models with linear feature aggregation mechanisms, such as RWKV and Mamba [4, 5], which emerge as efficient solutions for processing lengthy sequences while maintaining transformer-like capabilities. These architectures demonstrate that linear attention can effectively capture long-range dependencies without the quadratic computational overhead of traditional self-attention. The adaptation of RWKV architectures for diffusion models has shown promising results in addressing computational complexity limitations. Diffusion-RWKV [28] demonstrates that RWKV-based architectures can efficiently handle patchified inputs in diffusion models while significantly reducing spatial aggregation complexity, making them exceptionally adept at processing high-resolution images without requiring windowing or group cached operations. This work establishes the feasibility of linear attention mechanisms in generative modeling tasks, achieving performance comparable to transformer-based approaches while reducing computational overhead. VRWKV [8] adapts linear attention principles to vision tasks, demonstrating robust global processing capabilities while efficiently handling sparse inputs and maintaining linear computational complexity. The bidirectional processing capabilities of VRWKV’s Bi-WKV mechanism offer particular promise for tasks requiring comprehensive spatial-temporal understanding (e.g. video editing). Building upon the success of Diffusion-RWKV in image generation, the extension to video editing tasks represents a natural progression toward more efficient video processing architectures.
2.3 Computational Efficiency in Video Processing
The computational demands of video processing have motivated various efficiency-oriented approaches that directly relate to our linear attention integration [29, 30]. Current video editing methods exhibit significant variations in computational requirements, with training-free methods like TokenFlow [16], Rerender [31], and RAVE [32] showing superior efficiency while maintaining satisfactory performance. However, these methods still face scalability challenges when processing high-resolution videos or long sequences due to underlying quadratic attention mechanisms [23, 33, 34]. The hierarchical nature of traditional attention mechanisms, while beneficial for capturing multi-scale features, often results in increased parameter counts and computational overhead, particularly challenging for real-time video processing applications [35]. This limitation becomes especially pronounced in video editing scenarios where maintaining temporal consistency requires processing multiple frames simultaneously [14, 11, 15]. Our work addresses these computational limitations by proposing VRWKV-Editor, which leverages the linear complexity of VRWKV to achieve superior computational efficiency while maintaining high-quality video editing performance. Unlike previous approaches that rely on quadratic attention mechanisms, our method enables seamless processing of high-resolution videos without requiring window operations or hierarchical attention patterns.
3 Background
This section provides the necessary foundations that form the backbone of our work. We first discuss latent variable refinement techniques that improve efficiency while preserving quality. We then review diffusion models in the context of video generation and highlight their extension from image-based frameworks to spatio-temporal domains, followed by a detailed overview of the VRWKV architecture, which offers a linear-complexity alternative to traditional attention mechanisms. These components form the conceptual backbone for our proposed method.
3.1 Diffusion Models and Latent Variable Refinement
Diffusion models represent [36] a powerful class of generative models that operate through a progressive noise addition and removal process. Starting with an original image $x_{0}$, the forward diffusion process gradually corrupts the data by adding Gaussian noise over $T$ timesteps. At each step $t$, the noisy image $x_{t}$ is obtained through:
$x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t},$ (1)
where $\alpha_{t}$ represents a noise schedule parameter and $\epsilon_{t}\sim N(0,I)$ is Gaussian noise with the same dimensions as $x_{0}$.
The reverse process learns to denoise these corrupted images, effectively modeling the data distribution. For conditional generation tasks, we can extend this framework to model the conditional distribution $p_{\theta}(x_{t}|c,x_{0})$, where $c$ represents conditioning information such as text prompts or class labels. The conditional model learns:
$p_{\theta}(x_{0}|c)=\int p_{\theta}(x_{0:T}|c)dx_{1:T},$ (2)
$p_{\theta}(x_{0:T}|c)=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t},c),$ (3)
where $x_{T}\sim N(0,I)$ represents pure noise.
The model parameters $\theta$ are optimized by maximizing the variational lower bound of the log-likelihood:
$\mathcal{L}=\mathbb{E}_{x_{0},c,t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|^{2}\right],$ (4)
where the network $\epsilon_{\theta}$ learns to predict the noise added at each timestep given the noisy image $x_{t}$, timestep $t$, and conditioning information $c$.
To improve computational efficiency, latent diffusion models perform the diffusion process in a compressed latent space rather than directly in pixel space. This approach utilizes a pre-trained autoencoder consisting of an encoder $E$ that maps images to latent representations $z=E(x)$ and a decoder $D$ that reconstructs images from latents.
The diffusion process then operates in this latent space, learning the distribution of encoded representations. This approach significantly reduces memory and computational costs while maintaining high-quality generation capabilities [37]. During inference, the model generates latent representations through the reverse diffusion process, which are then decoded back to pixel space using the pre-trained decoder.
This framework enables efficient conditional image generation by combining the expressiveness of diffusion models with the computational advantages of latent space operations, making it particularly suitable for high-resolution image synthesis tasks.
3.2 Diffusion Models for Video Processing
Beyond their success in image generation, diffusion models have recently been extended to video synthesis, where they enable high-quality spatiotemporal generation and editing [38, 39, 40]. These models operate by learning to reverse a gradual noise addition process, where data is progressively corrupted through a forward Markov chain and then restored through a learned reverse process.
The transition from image to video generation involves inflating 2D models to handle spatio-temporal dimensions, typically through pseudo-3D convolutions where 3×3 kernels are replaced by 1×3×3 kernels, and temporal attention layers are added to transformer blocks [22, 41]. However, this extension introduces significant computational challenges due to the quadratic complexity of attention mechanisms when processing multiple frames simultaneously.
3.3 VRWKV Architecture
Figure 2: VRWKV [8] module employed on VRWKV-Editor. The VRWKV architecture includes identical VRWKV encoder layer, an average pooling layer, and a linear prediction head. Q-Shift denotes the quad-directional shift method tailed for vision tasks.
VRWKV consists of three main components: Receptance $(R)$, Key $(K)$, and Value $(V)$ transformations, combined with a time-mixing mechanism specifically adapted for vision tasks. For a given input sequence of video frames $X_{t}\in\mathbb{R}^{f\times c\times h\times w}$, where $f$ is the number of frames, $c$ is the number of channels, and $(h,w)$ are the spatial dimensions, the VRWKV operations are defined as:
$R_{t}=W_{R}\otimes(\mu_{R}\odot X_{t}+(1-\mu_{R})\odot X_{t-1}),$ (5)
$K_{t}=W_{K}\otimes(\mu_{K}\odot X_{t}+(1-\mu_{K})\odot X_{t-1}),$ (6)
$V_{t}=W_{V}\otimes(\mu_{V}\odot X_{t}+(1-\mu_{V})\odot X_{t-1}).$ (7)
$W_{R}$, $W_{K}$, $W_{V}$ are learnable projection matrices, $\mu_{R}$, $\mu_{K}$, $\mu_{V}$ are learnable interpolation factors, and $\otimes$ denotes a spatiotemporal convolution, defined as a standard $2D$ spatial convolution applied on a temporally interpolated frame representation $\mu\odot X_{t}+(1-\mu)\odot X_{t-1}$. This differs from standard 3D convolution since it only considers the current and previous frame, rather than a full temporal kernel.
The core innovation of VRWKV lies in its weighted key-value mechanism that maintains temporal coherence:
$wkv_{t}=\operatorname{Bi-WKV}(K,V)_{t}=\frac{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}+e^{u+k_{t}}}.$ (8)
where $T$ represents the total number of tokens, $w$ and $u$ are two $d$-dimensional learnable vectors that represent channel-wise spatial decay and the bonus indicating the current token, respectively. $k_{t}$ and $v_{t}$ denotes $t$-th feature of $K$ and $V$.
The final output incorporates both the weighted key-value information and the receptance gate:
$O_{t}=\sigma(R_{t})\odot{wkv}_{t},$ (9)
where $\sigma$ represents the sigmoid activation function.
The output $O_{t}$ from the time-mixing block is then processed through a channel-mixing block that enhances feature representation across different channels. The channel-mixing mechanism operates as follows:
$R_{c,t}$ $=W_{R_{c}}\otimes(\mu_{R_{c}}\odot O_{t}+(1-\mu_{R_{c}})\odot O_{t-1}),$ (10)
$K_{c,t}$ $=W_{K_{c}}\otimes(\mu_{K_{c}}\odot O_{t}+(1-\mu_{K_{c}})\odot O_{t-1}),$ (11)
where $W_{R_{c}}$ and $W_{K_{c}}$ are channel-mixing projection matrices, and $\mu_{R_{c}}$, $\mu_{K_{c}}$ are learnable interpolation parameters for the channel-mixing component. The channel-mixed output is computed as:
$CM_{t}=\sigma(R_{c,t})\odot\max(K_{c,t}\cdot W_{c},0)^{2},$ (12)
where $W_{c}$ is a learnable transformation matrix and the squared ReLU activation $\max(\cdot,0)^{2}$ provides enhanced non-linearity to model channel interactions. The final layer output combines both time-mixing and channel-mixing components through a residual connection:
$Y_{t}=X_{t}+O_{t}+CM_{t},$ (13)
Figure 3: Pipeline of VRWKV-Editor: Given a text–video pair (e.g., “A man with a backpack hikes on a rocky terrain”) as input, our method leverages pretrained text-to-image diffusion models for text-to-video generation. The input video is first encoded into a discrete latent space, after which our U-Net architecture predicts the injected noise. During inference, a novel video is synthesized by inverting the discrete noise from the input video, guided by an edited prompt (e.g., “An astronaut with a jetpack floats above a Martian landscape, with red rocky terrains and tall”).
This dual-mixing architecture enables VRWKV to capture both temporal dependencies through the time-mixing block and cross-channel feature interactions through the channel-mixing block, providing comprehensive spatiotemporal modeling capabilities for vision tasks.
4 Method
This section outlines our method, we begin by formulating the problem and highlighting the computational efficiency challenges encountered in existing architectures. This serves as the primary motivation for our approach, we then introduce the proposed architecture.
4.1 Problem Formulation
Given a video sequence $V=\{v_{i}\}_{1}^{n}$ with $n$ frames, and a text prompt $P$ describing $V$, our goal is to generate an edited video $V^{{}^{\prime}}=\{v_{i}^{{}^{\prime}}\}_{1}^{n}$ that aligns with new prompts $P^{{}^{\prime}}$ provided by the user, where $v_{i}^{{}^{\prime}}\in\mathbb{R}^{C\times H\times W}$ is a frame with C channels, height H and width W.
4.2 Computational Efficiency Challenges
The computational demands of video processing, particularly for high-resolution content and long sequences, present a fundamental bottleneck in current methods [38, 42]. Traditional ViTs process high-resolution frames resulting in quadratic growth of memory and computation due to the pairwise token interactions in the self-attention mechanism. For example, a $\times2048$ frame (16,384 tokens with a patch size of 16) requires dramatically more memory and processing time than lower-resolution content, making real-time applications impractical. The classical self-attention is computed as follows:
$\text{Attn}(Q,K,V)=\text{softmax}(QK^{\top})V,$ (14)
where Q, K and V represents the Querie, Key and Value respectively. Although effective in capturing long-range dependencies, this operation scales poorly with sequence length.
Variants such as AFT [43] replace the dot-product operation in standard attention with element-wise multiplication, leading to a linearized formulation that avoids explicit query–key interactions. Specifically, AFT introduces learnable positional biases $w_{t,i}$ into the exponential weights:
$\text{Attn}^{+}(W,K,V)_{t}=\frac{\sum_{i=1}^{t}e^{w_{t,i}+k_{i}}\odot v_{i}}{\sum_{i=1}^{t}e^{w_{t,i}+k_{i}}},$ (15)
where the terms $w_{t,i}$ encode contextual dependencies and act as learnable position encodings. This design captures structured interactions across tokens while reducing the quadratic cost of standard attention. Nevertheless, even with such improvements, many methods still resort to hierarchical or windowed attention to further reduce memory and computation, at the expense of the global receptive field, which is crucial for modeling long-range dependencies in video understanding. Linear vision transformers, such as VRWKV, address these challenges by reformulating token interactions to achieve linear complexity with respect to sequence length. In RWKV, attention weights are expressed as a channel-wise time decay:
$w_{t,i}=-(t-i)w,\quad w\in\mathbb{R}_{\geq 0}^{d},$ (16)
[…]
4.3 Overview of VRWKV-Editor
Tune-A-Video [11] reduced the high computational complexity of spatio-temporal attention (ST-Attn) in generating videos with increasing frames by introducing a novel sparse version of causal attention mechanism, where the attention matrix is computed between frame $z_{v_{i}}$ and two previous frames $z_{v_{1}}$ and $z_{v_{i-1}}$, reducing the complexity from $O((dT)^{2})$ to $O(2dT^{2})$. However, the quadratic complexity still remains a bottleneck for long sequences of video frames, limiting their ability to efficiently process high-resolution images and lengthy sequences, posing a significant barrier to their broader application. To overcome this issue, we propose a linear attention mechanism based on the VRWKV (Vision Receptance Weighted Key Value) framework, where we replace the traditional quadratic attention computation with a receptance-weighted key-value operation that scales linearly with sequence length, reducing the complexity from $O(2dT^{2})$ to $O(2dT)$. Our approach leverages the VRWKV’s inherent ability to capture long-range dependencies through its recurrent formulation while maintaining spatial awareness through learnable shift operations and channel mixing.
An overview of our approach is depicted in Figure 3. Given a text-video pair as input, our model leverages the pretrained T2I diffusion models for T2V generation by encoding the input video into a discrete space, then we predict the added noise using our U-Net architecture enhanced with VRWKV modules. During inference, we sample a novel video from the discrete noise inverted from the input video, guided by an edited prompt.
4.4 Vision-Recurrent Weighted Key-Value (VRWKV) Module
During our experiments, we observed that traditional CNN-based spatiotemporal modules failed to capture complex spatiotemporal dependencies, while adding more cross-attentions or Temporal Attentions only increased the complexity of the model without improving the quality of edited videos or maintaining their consistency. Inspired by the RWKV architecture’s ability to combine the strengths of RNNs and Transformers, we introduce a variant of Vision-Recurrent Weighted Key-Value (VRWKV) module specifically designed for video spatiotemporal modeling. The VRWKV module addresses key limitations of traditional attention mechanisms in video editing. Unlike quadratic attention mechanisms, VRWKV operates with linear complexity $O(d)$, making it more efficient for processing long video sequences (Table 2). The module maintains a recurrent state that captures temporal dependencies across frames while preserving spatial relationships. Moreover, the weighted key-value mechanism allows for dynamic focus on relevant spatiotemporal features.
Figure 4: Qualitative comparison between VRWKV-Editor [44] and CAMEL [45], CCEdit [46], ControlVideo [20], Swin-Editor [44] and Vid2Vid-zero [10]. The first row presents the input video. The figure highlights two different scenarios: Object editing and object-background editing . The results showcase the superiority of our framework in editing videos across various scenarios, producing consistent and high quality outputs.
5 Experiments
We conduct experiments to test the performance of VRWKV-Editor with existing video-editing methods. Our evaluation encompasses computational efficiency metrics, generation quality assessments, and temporal coherence analysis across various video editing tasks.
Figure 5: Comparison of runtime between our VRWKV Sparse Attention (as deployed in our model) and the Sparse Causal Attention used in Tune-A-Video [11] is presented, showing forward pass, backward pass, and total execution times across varying input sizes. The execution time for VRWKV Sparse Attention scales linearly with input size, whereas the Sparse Causal Attention in Tune-A-Video exhibits quadratic growth.
5.1 Implementation Details
For VRWKV, we implemented a single layer of RWKV-V4 (Figure 3) with local initialization mode, where the spatial mixing parameters are initialized to 1.0, making it behave more like local attention initially. The number of attention heads is set to 16 with 88 dimension for each head (total embedding dimension of 1408). The vision-specific hyperparameters include patch size of 16×16, input frame resolution of 512×512, spatial shift pixel of 1, and channel gamma of 1/4 for directional shifting across four spatial directions (up, down, left, right), enabling efficient processing of high-resolution frames without windowing operations.
During inference, we implemented the DDIM sampler [47] alongside classifier-free guidance [48] of $7.5$, controlling how strongly the model follows the text prompt. All experiments are conducted on a single machine with an NVIDIA A100-SXM4-80GB GPU.
Evaluated Models: We compared our method against six recent baselines: CAMEL [45], CCEdit [46], Tune-A-Video [11], Video-P2P [14], Vid2Vid-zero [10] and Swin-Editor [44]. Experiments are conducted with their official codes and configurations.
5.2 Computational Efficiency Analysis:
Timing Performance Comparison: Our timing experiments reveal significant computational advantages of VRWKV-Editor over window-based attention architectures. Figure 6 presents the forward and backward pass timing results for different input sequence lengths. The results demonstrate that VRWKV-Editor achieves substantial speedups that increase with input size, confirming the theoretical $O(n)$ vs $O(n^{2})$ complexity advantage. For large inputs (4096×4096), our method achieves up to 3.68× speedup in total computation time.
Memory Efficiency Analysis: VRWKV-Editor exhibits significantly better memory scaling compared to traditional attention-based architectures. While memory usage of conventional attention mechanisms grows quadratically with sequence length due to attention matrix storage, VRWKV-Editor maintains linear memory growth. At 4K resolution, VRWKV-Editor uses approximately 60% less GPU memory than swin attention based approaches.
Quantitative Results: In Table 1, we report the quantitative results of VRWKV-Editor versus state-of-the-art video editing methods. As the results show in Figures 1, 3 and Table 1, our model achieves a trade-off between temporal consistency, text alignment (for objects, background, and style) and high-quality video editing while significantly reducing computational complexity. This is also evidenced by the performance obtained in both automatic evaluations and user studies.
Qualitative Evaluation: Figure 3 showcases a visual comparison between our method and the other baselines. Further results are shown in Figures 7 and 8. We conducted our evaluation by editing objects, background, style or combination between them.
Table 1: Quantitative comparison between our method and evaluated baselines in terms of frame consistency, textual alignment, and runtime (minutes).
Method Frame Consistency Textual Alignment Runtime [min]
Clip Score $\uparrow$ User Vote $\uparrow$ Pick Score $\uparrow$ User Vote $\uparrow$ Training $\downarrow$ Inference $\downarrow$
CAMEL 93.38 14.20 26.67 8.60 13.00 1.50
CCEdit 93.07 10.10 25.70 11.20 14.50 2.00
Tune-A-Video 95.22 17.30 28.18 20.10 11.00 0.50
Video-P2P 96.40 11.60 25.82 8.10 22.00 2.00
Vid2Vid-zero 94.18 8.60 26.98 14.00 3.00 3.00
Swin-Editor 95.47 18.40 28.05 18.70 10.00 0.30
VRWKV-Editor 95.75 19.80 27.90 19.30 8.00 0.16
Figure 6: Computational timing comparison between VRWKV Attention and Swin Attention deployed in our model showing forward pass, backward pass, and total execution times across different input sizes. VRWKV-Editor demonstrates increasing efficiency gains with larger inputs, achieving up to 3.68× speedup for 4096×4096 inputs.
Table 2: Performance according to Video Length.
Video Length Swin-Attn Time (s) VRWKV-Attn Time (s) Memory Ratio
16 frames 2.01 1.02 1.97×
32 frames 3.99 1.78 2.24×
64 frames 7.95 3.54 2.24×
128 frames 15.89 7.05 2.25×
6 Limitations:
A key conceptual limitation of the current VRWKV-Editor lies in its use of VRWKV in a Transformer-style mode, without returning recurrent states across time. While this formulation aligns with conventional transformer architectures, it does not leverage VRWKV’s native RNN mode, which enables true sequential processing with memory.
Adapting VRWKV-Editor to operate in its RNN mode presents a promising research direction. By returning and propagating states during inference, the model could process arbitrarily long videos chunk by chunk while maintaining temporal coherence through memory linking. This would exploit the fundamental advantage of RNNs over transformers, which require the full sequence as input or resort to independent chunking, thereby limiting scalability. Such an extension could unlock VRWKV’s potential for efficient editing of very large-scale videos.
7 Conclusion:
In this paper, we presented VRWKV-Editor, a novel video editing framework that leverages VRWKV’s linear spatio-temporal attention mechanism to address the computational limitations of traditional quadratic attention methods. The exploration of VRWKV in video editing task demonstrates its potential to be a viable alternative to traditional video understanding models, successfully reducing computational complexity while maintaining high-quality editing performance. The successful integration of RWKV architectures into video editing workflows opens new possibilities for efficient spatiotemporal modeling and real-time video processing applications. Future work should focus on optimizing low-level GPU kernels and extending the framework to broader video understanding tasks. This research contributes to the advancement of efficient video processing architectures and demonstrates the potential of linear attention mechanisms in computer vision tasks.